What is the mass of (4.9x10^24) formula units of sodium chloride?


1
Expert's answer
2021-07-19T04:04:29-0400

Number of molecules

= number of molecules/ 6.02 × 1023 .


Number of moles = 4.9 × 1024/ 6.02 × 1023.


= 8.13 moles .


Mass of NaCl = ( 8.13 × 58.5 ) g = 476.16 g .